OpenStudy (anonymous):

Israel added $120 to his savings. This added amount represents 1/6 of his total savings. If s represents the total savings, which equation could be used to determine the value of s? A. 120s=1/6 B. 1/6s=120 C. S=120+1/6 D. 120+s=1/6

OpenStudy (anonymous):

im not good with story probs but i can try

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

im thinking A or B

OpenStudy (anonymous):

which is the best answer?

OpenStudy (anonymous):

im thinking A

OpenStudy (anonymous):

like i said im not good with story probs lol

OpenStudy (anonymous):

The sum of three consecutive integers is 78 . Determine the smallest integer.

OpenStudy (anonymous):

what about this ^

OpenStudy (anonymous):

was A right

OpenStudy (anonymous):

@ilovehim14<3

OpenStudy (anonymous):

@Vallarylynn16

OpenStudy (anonymous):

let me yahoo lol

OpenStudy (anonymous):

idk, im not done , The sum of three consecutive integers is 78 . Determine the smallest integer.

OpenStudy (anonymous):

Let the first integer be x. Then the second is x+1 and the third is x+2. x+x+1+x+2 = 228 3x + 3 = 228 3x = 225 x = 225/3 x = 75 Therefore the integers are 75, 76 and 77.

OpenStudy (anonymous):

@Cpt.Sparz , @AuroraB , @Vallarylynn16

OpenStudy (anonymous):

Sherrie receives 5 times as many phone calls as Carrie, and Carrie receives 5 less calls than Mary. If the total number of combined calls for them was 68, how many calls did Carrie receive? A. 9 calls B. 14 calls C. 19 calls D. 45 calls

OpenStudy (anonymous):

@Cpt.Sparz

OpenStudy (anonymous):

i cant sorry

OpenStudy (anonymous):

Yessir, I think it's 14...but I could be wrong..